The amount of medicine in micrograms, A, in a persons bloodstream t minutes after a certain quantity of it has been inhaled can be approximated by the equation A=-4t^2 + 40t. How long after inhalation will there be 84 micrograms of Medicine in the bloodstream?
Please help if someone can!

Answer by ankor@dixie-net.com(22740)

The amount of medicine in micrograms, A, in a persons bloodstream t minutes after a certain quantity of it has been inhaled can be approximated by the equation A=-4t^2 + 40t.
How long after inhalation will there be 84 micrograms of Medicine in the bloodstream?
:
A = 84
so we have
4t^2 + 40t = 84
simplify, divide by 4
t^2 + 10t = 21
a quadratic equation
t^2 + 10t - 21 = 0
Use the quadratic formula to find t; a=1, b=10, c=-21
I got a positive solution of 1.78 min